# Xingzhi Zhou1,3, Xin Dong2, Chunhao Li1, Yuning Bai4, Yulong Xu5, Ka Chun Cheung3, Simon See6, Xinpeng Song2 ✉Runshun Zhang3, ✉Xuezhong Zhou2, and ✉Nevin L. Zhang1

### v1

@@ Abstract
Traditional Chinese medicine (TCM) relies on specific combinations of herbs in prescriptions to treat symptoms and signs, a practice that spans thousands of years. Predicting TCM prescriptions presents a fascinating technical challenge with practical implications. However, this task faces limitations due to the scarcity of high-quality clinical datasets and the intricate relationship between symptoms and herbs. To address these issues, we introduce DigestDS, a new dataset containing practical medical records from experienced experts in digestive system diseases. We also propose a method, TCM-FTP (TCM Fine-Tuning Pre-trained), to leverage pre-trained large language models (LLMs) through supervised fine-tuning on DigestDS. Additionally, we enhance computational efficiency using a low-rank adaptation technique. TCM-FTP also incorporates data augmentation by permuting herbs within prescriptions, capitalizing on their order-agnostic properties. Impressively, TCM-FTP achieves an F1-score of 0.8031, surpassing previous methods significantly. Furthermore, it demonstrates remarkable accuracy in dosage prediction, achieving a normalized mean square error of 0.0604. In contrast, LLMs without fine-tuning perform poorly. Although LLMs have shown capabilities on a wide range of tasks, this work illustrates the importance of fine-tuning for TCM prescription prediction, and we have proposed an effective way to do that.
@@ I INTRODUCTION
Traditional Chinese Medicine (TCM) has been an indispensable part of healthcare for the Chinese population for thousands of years. TCM employs a wide range of practices, including herbal medicine, acupuncture, cupping therapy, and tuina massage [1]. Herbal medicine is the primary treatment modality of TCM. It has been shown to effectively treat the novel coronavirus (COVID-19), resulting in improved cure rates and reduced mortality [2, 3].
@@
Herbal prescriptions require doctors to assess patient symptoms using the four diagnostic methods: observation (wang), listening and smelling (wen), questioning (wen), and pulse-taking (qie), guided by the principle of li-fa-fang-yao [4]. Training a doctor involves extensive experience and continuous practical feedback, resulting in a prolonged training period. This contributes to a shortage of TCM doctors and limited TCM resources. Developing an effective prescription predictor is an approach to alleviate challenges such as doctor shortages and the need for prolonged training periods.
Prescription prediction in TCM involves designing a computational system capable of predicting the appropriate prescription based on given symptoms. This system models the complex relationships between symptoms and herbs [5]. Language generation techniques show particular promise in TCM prescription prediction, treating prescription generation as a machine translation problem and solving it using a sequence-to-sequence (seq2seq) model [6]. After numerous attempts by various researchers [7, 8, 9, 10], generative models have shown positive results in TCM prescription prediction. However, current models are beset with the following limitations:
  • Scarcity of clinical datasets. Existing works often rely on datasets derived from classical documents rather than high-quality datasets from clinical records, which introduces noisy information. This results in models that cannot effectively provide personalized prescription recommendations [11]. Moreover, there is currently a lack of high-quality clinical data for prescription prediction [12].
  • Sub-optimal Performance. Existing models frequently yield results that do not meet expectations and may even mislead users due to their sub-optimal prediction accuracy [12]. However, relying solely on clinical data is insufficient to enhance model performance [13]. It is necessary to integrate more advanced language model techniques to improve the representation and encoding capabilities of prescription prediction models.
  • Lack of herb dosage. Existing models overlook the crucial prediction of dosage weights, a critical component in TCM for effective disease treatment [5, 14]. Additionally, there is a lack of evaluation metrics specifically for herb dosage prediction.
@@
Inspired by the robust predictive capabilities of large language models (LLMs), these challenges motivate us to develop a high-quality prescription dataset and leverage LLMs to construct an advanced prescription predictor. In this study, we introduce a high-quality prescription dataset DigestDS derived from clinical medical records and propose TCM-FTP (Fig. 1), a novel TCM prescription generation model based on LLMs. Our goal is to leverage the capabilities of LLMs to overcome the limitations of current methodologies. Specifically, we utilize a low-rank adaptation technique (LoRA) [15] for efficient LLM fine-tuning. To take advantage of the order-agnostic nature of herb prescriptions, we implement data augmentation by randomizing the sequence of herbs in the training data. We validated our model on DigestDS, achieving a Precision of 0.7951, a Recall of 0.8113, and an F1-score of 0.8031. This represents a significant improvement over the best performance achieved by previous methods.
Our main contributions are outlined as follows:
@@
Figure 1: Workflow of the TCM-FTP. Our work consists of four parts: ”Data Collection” involves gathering and organizing raw data; ”Data Processing” includes data preprocessing, prompt design, and integrating data augmentation to create a fine-tuning dataset; ”Fine-tuning” utilizes the ShenNong LLM and LoRA technique to optimize the model; and ”Evaluation” assesses the outcomes using both quantitative and qualitative evaluation metrics.
@@
  • We construct a TCM clinical diagnostic dataset DigestDS derived from practical records in real-world TCM clinical scenarios. This dataset serves as a foundational resource for understanding the nuanced relationships between symptoms and effective prescriptions in TCM.
@@
  • We introduce TCM-FTP, a novel LLM-driven method for generating TCM prescriptions. Our model excels at capturing the intricate interplay between symptoms and herbal prescriptions, marking a significant advancement in TCM prescription prediction. Additionally, we have designed the NMSE evaluation metric for dosage prediction. To the best of our knowledge, our work is the first in the prescription prediction field to design dosage prediction metrics tailored for real-world scenarios.
@@
  • We empirically validate TCM-FTP against existing models on our dataset, showcasing its superior performance in predicting herbs and dosages. Additionally, we conduct a thorough analysis of factors such as the impact of foundational models, learning rates, the number of herb permutations, and the decoding parameters in inference. Expert case studies by TCM professionals further provide valuable insights into the adaptability and robustness of our innovative approach.
@@ II-B TCM Herbal Prescription Prediction
Research on TCM prescription prediction mainly falls into three categories: topic model-based, graph model-based, and language model-based approaches. Topic model-based approaches treat relationships between symptoms and herbs as that of documents and topics [16, 17, 18]. These approaches rely on statistical relationships, lacking a semantic understanding of symptoms. Graph-based approaches construct a medical knowledge graph to model the relationships between symptoms and herbs [19, 14, 20]. However, these approaches also lack consecutive semantic information regarding symptoms. Language model-based approaches are a more promising way to model the complicated relationships between symptoms and herbs.
Language model-based prescription prediction models take patient symptom descriptions as input and generate herbal prescriptions sequentially. TCM Translator [21] uses transformer architectures to distill context vectors from symptoms and LSTM [22] as the decoder. AttentiveHerb [7] employs a seq2seq [6] model with dual attention mechanisms to distinguish primary from secondary symptoms and map herb-symptom interactions using clinical data. Herb-Know [8] utilizes herb descriptions to model associations with symptoms and evaluates whether herb effects align with symptom descriptions. TCMBERT [9] integrates transfer learning by initially training an ALBERT [23] on TCM-related documents and fine-tuning an LSTM-based seq2seq model for context vector extraction. RoKEPG [10] incorporates additional herb knowledge for fine-tuning prescription prediction models.
@@ II-C Large Language Models in TCM
LLMs have made significant strides in various NLP tasks. Notable examples like ChatGPT and GPT-4.0 [24] have attracted considerable attention, although specifics about their architectures, parameters, and training strategies from OpenAI remain undisclosed. The emergence of open-sourced LLMs has sparked widespread interest among researchers. Prominent open-sourced models include LLaMA [25], Bloom [26], and ChatGLM [27], among others. While LLMs excel in standard NLP tasks, they often struggle with specialized domains requiring specific knowledge, such as medicine or finance. Supervised fine-tuning has become a standard approach to enhance LLMs for these domains by incorporating specialized knowledge.
In the context of TCM, several innovative models have been proposed. Bentsao [28] utilizes supervised fine-tuning on LLaMA, integrating structured and unstructured knowledge from CMeKG [29]. Huatuo [30] leverages both original and distilled data from ChatGPT, incorporating a Reinforcement Learning from Artificial Intelligence Feedback (RLAIF) mechanism and employing Proximal Policy Optimization (PPO) during fine-tuning. Zhongjing [31] applies continual pretraining to inject domain knowledge and uses supervised fine-tuning on a Chinese multi-turn medical dialogue dataset, complemented by Reinforcement Learning from Human Feedback (RLHF). ShenNong [32] builds on LLaMA with LoRA-based fine-tuning on an instructional dataset derived from ChatGPT and a traditional Chinese medicine knowledge graph, benefiting from a large-scale dataset of over 110,000 instructions.
@@ II-D Parameter Efficient Fine-Tuning
Parameter Efficient Fine-Tuning (PEFT) utilizes a small amount of parameters to fine-tune a large language model effectively. Assuming there is a pretrained model $f_{\theta}(y|x)$, PEFT seeks to adjust a limited number of parameters, $\Delta\theta$, such that $|\theta|\gg|\Delta\theta|$. In contrast to conventional fine-tuning, which updates all parameters, denoted as $|\theta|=|\Delta\theta|$, and requires significant computational resources, PEFT selectively updates a limited number of learnable parameters to achieve results comparable to complete fine-tuning.
PEFT encompasses three primary methodologies: adapter [33], p-tuning [34, 35, 36], and LoRA [15]. Adapter tuning [33] introduces PEFT by appending additional learnable modules, named adapters, to every layer of the pretrained model, exclusively fine-tuning these adapters while maintaining the original parameters. In contrast, p-tuning leveraged prompt engineering techniques, fine-tuning learnable vectors as new tokens and inserting these tokens into the multi-head self-attention layer (MSA) [34, 35, 36]. In this work, we employ LoRA [15] to efficiently fine-tune LLMs on the target dataset.
@@ III-A1 Data collection
We collect outpatient medical record data generated by specialists in TCM hospital  [^1] Dr. Runshun Zhang, Dr. Yuning Bai, etc. over a span from 2008 to 2022. The prescriptions specifically focus on digestion system disorders.
@@ III-B TCM-FTP
Given a prescription dataset $\mathcal{P}_{train}$ with $\{s,\{h_{i},w_{i}\}^{i\in[k]}\}$, our primary goal is to use a language model to maximize the estimation of conditional probability:
$\max\limits_{\theta}\mathop{\mathbb{E}}\limits_{s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}}\sum\limits_{i}^{k}\log P((h_{i},w_{i})|h_{<i},w_{<i};s;\theta),$ (1)
@@
To leverage the auto-regressive mechanism in LLMs, we concatenate the herbs and dosages into one string. We denote concatenated string $\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]})$ as $y$. Consequently, our objective function becomes:
$\max\limits_{\theta}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}\\ y=\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]})\end{subarray}}\sum\limits_{j=1}^{|y|}\log P(y_{j}|\{y_{<j}\};s;\theta),$ (2)
where $\{y_{<j}\}$ denotes the set of the tokens from index $1$ to index $j-1$ in $y$, and $|y|$ represents the number of tokens in $y$.
@@ III-B1 Low-rank Adaptation
We employ the LoRA [15] for PEFT to conserve computational resources. Beginning with a pre-trained LLM $f_{\theta}$, we encounter two primary types of parametric functions within our model framework: linear and embedding functions. These functions are mathematically described as:
${\mathcal{F}_{\text{linear}}}(x)$ $={W_{l}}\cdot x,$ (3)
${\mathcal{F}_{\text{emb}}}(x)$ $=\text{EMB}\left(x;{W_{e}}\right),$ (4)
where EMB$(x;\cdot)$ represents the embedding operator that selects the $x$-th column from the specified matrix, and $W_{l}$ and $W_{e}$ denotes the parameters for the linear and embedding functions, respectively. For fine-tuning, we introduce updates via low-rank matrices $A$ and $B$, leading to modifications in the original functions:
${\mathcal{F}_{\text{linear}}}{(x)^{\prime}}$ $={\mathcal{F}_{\text{linear}}}(x)+{A_{l}}\cdot{B_{l}}\cdot x,$ (5)
${\mathcal{F}_{\text{emb}}}{(x)^{\prime}}$ $={\mathcal{F}_{\text{emb}}}(x)+{A_{e}}\cdot\text{EMB}\left(x;{B_{e}}\right),$ (6)
where $A_{l}$, $B_{l}$, $A_{e}$, and $B_{e}$ are the learnable parameters associated with the linear and embedding functions. With the updates to the low-rank matrices represented by $\Delta\theta$, our goal is to maximize the expected conditional probability:
$\max\limits_{\Delta\theta}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}\\ y=\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]}\end{subarray}}\sum\limits_{j=1}^{|y|}\log P(y_{j}|\{y_{<j}\};s;\theta+\Delta\theta).$ (7)
@@ III-C Baselines
To demonstrate the performance of our model, we compare it with the following baselines:
(1) Topic Models and Multi-label Classification Models:
  • LinkLDA [37] uses Link Latent Dirichlet Allocation [38] to model relationships between symptoms and corresponding herbs.
  • LinkPLSALDA [39] combines LDA [40] and PLSA [41] for topic modeling.
  • MLKNN [42] is an enhanced K-nearest neighbor (KNN) approach for multi-label classification (MLC).
(2) TCM Prescription Prediction Models:
  • PTM [5] excels in recommending herbs based on symptoms and predicting symptoms from herbs.
  • TCMPR [43] utilizes sub-networks for symptom description and information extraction.
  • KDHR [44] integrates herb properties and features using multi-level GCN layers.
  • PresRecST [12] follows a systematic approach with three stages for clinical predictions in TCM.
(3) Pre-trained Language Models:
  • Mengzi (T5-base) [45] is a T5-based [46] seq2seq model pre-trained on Chinese text and fine-tuned on the specific dataset as a competitive baseline.
  • GPT-3.5 and GPT-4.0 [24]: We also include the pre-trained LLMs without fine-tuning as baselines.

### v2

# TCM-FTP: Fine-Tuning Large Language Models for Herbal Prescription Prediction

@@ Abstract
Traditional Chinese medicine (TCM) has relied on specific combinations of herbs in prescriptions to treat various symptoms and signs for thousands of years. Predicting TCM prescriptions poses a fascinating technical challenge with significant practical implications. However, this task faces limitations due to the scarcity of high-quality clinical datasets and the complex relationship between symptoms and herbs. To address these issues, we introduce DigestDS, a novel dataset comprising practical medical records from experienced experts in digestive system diseases. We also propose a method, TCM-FTP (TCM Fine-Tuning Pre-trained), to leverage pre-trained large language models (LLMs) via supervised fine-tuning on DigestDS. Additionally, we enhance computational efficiency using a low-rank adaptation technique. Moreover, TCM-FTP incorporates data augmentation by permuting herbs within prescriptions, exploiting their order-agnostic nature. Impressively, TCM-FTP achieves an F1-score of 0.8031, significantly outperforming previous methods. Furthermore, it demonstrates remarkable accuracy in dosage prediction, achieving a normalized mean square error of 0.0604. In contrast, LLMs without fine-tuning exhibit poor performance. Although LLMs have demonstrated wide-ranging capabilities, our work underscores the necessity of fine-tuning for TCM prescription prediction and presents an effective way to accomplish this.
@@ I INTRODUCTION
Traditional Chinese Medicine (TCM), integral to Chinese healthcare for millennia, encompasses practices such as herbal medicine, acupuncture, cupping therapy, and tuina massage. As the primary modality of TCM, herbal medicine has demonstrated effectiveness in treating COVID-19, leading to improved cure rates and reduced mortality [1]. Herbal prescriptions in TCM require doctors to assess patient symptoms using four diagnostic methods—observation (wang), listening and smelling (wen), questioning (wen), and pulse-taking (qie)—guided by the principle of li-fa-fang-yao [2]. The extensive experience and continuous practical feedback required for training result in prolonged training periods, contributing to a shortage of practitioners and limited TCM resources. Developing an effective prescription predictor could alleviate challenges such as doctor shortages and lengthy training requirements.
Computational systems for prescription prediction in TCM model the complex relationships between symptoms and herbs [3]. Language generation techniques, treating prescription generation as a machine translation problem using sequence-to-sequence (seq2seq) models [4], have shown promise. Despite positive results from generative models [5, 6, 7, 8], current models still face the following limitations: (1) Scarcity of clinical datasets. Current models often rely on datasets derived from classical documents instead of high-quality clinical records, introducing noise and limiting the effectiveness of personalized prescription recommendations [9]. Moreover, there is a significant lack of high-quality clinical data for prescription prediction [10]. (2) Sub-optimal Performance. Existing models often underperform and may mislead users due to low prediction accuracy [10]. Relying solely on clinical data is insufficient to enhance model performance [11]; integrating advanced language modeling techniques is necessary to improve the representation and encoding capabilities of prescription prediction models. (3) Lack of herb dosage. Existing models overlook dosage weights, a critical component in TCM for effective treatment [3]. Furthermore, there is a lack of evaluation metrics specifically designed for herb dosage prediction.
@@
Motivated by the robust predictive capabilities of large language models (LLMs), we develop a high-quality prescription dataset, DigestDS, derived from clinical medical records, and propose TCM-FTP —a novel LLM-based TCM prescription generation model (Fig. 1). Specifically, we employ a low-rank adaptation technique (LoRA) [12] for efficient fine-tuning of the LLM. Exploiting the order-agnostic nature of herb prescriptions, we implement data augmentation by randomizing the sequence of herbs in the training data. Validated on DigestDS, our model achieves a precision of 0.7951, a recall of 0.8113, and an F1-score of 0.8031, significantly outperforming previous methods. Our main contributions are outlined as follows:
@@
Figure 1: Workflow of the TCM-FTP. Our work consists of four parts: “Data Collection” involves gathering and organizing raw data; “Data Processing” includes data preprocessing, prompt design, and integrating data augmentation to create a fine-tuning dataset; “Fine-tuning” utilizes the ShenNong LLM and LoRA technique to optimize the model; and “Evaluation” assesses the outcomes using both quantitative and qualitative evaluation metrics.
@@
  • We construct DigestDS, a TCM clinical diagnostic dataset derived from real-world clinical records, serving as a foundational resource for understanding the nuanced relationships between symptoms and effective prescriptions in TCM.
@@
  • We introduce TCM-FTP, a novel LLM-driven method for generating TCM prescriptions that excels at capturing the intricate interplay between symptoms and herbal treatments, marking a significant advancement in prescription prediction. Additionally, we introduce the Normalized Mean Square Error (NMSE) as an evaluation metric for dosage prediction, pioneering the first dosage prediction metric tailored to real-world scenarios in this field.
@@
  • We empirically validate TCM-FTP against existing models on our dataset, demonstrating its superior performance in predicting herbs and dosages. Expert case studies by TCM professionals further provide valuable insights into the adaptability and robustness of our innovative approach.
@@ II-B TCM Herbal Prescription Prediction
Research on TCM prescription prediction mainly falls into three categories: topic model-based, graph model-based, and language model-based approaches. Topic model-based approaches treat relationships between symptoms and herbs as that of documents and topics [13]. These approaches rely on statistical relationships, lacking a semantic understanding of symptoms. Graph-based approaches construct a medical knowledge graph to model the relationships between symptoms and herbs [14]. However, these approaches also lack consecutive semantic information regarding symptoms. Language model-based prescription prediction models take patient symptom descriptions as input and generate herbal prescriptions sequentially. TCM Translator [15] uses transformer architectures to distill context vectors from symptoms and LSTM as the decoder. AttentiveHerb [5] employs a seq2seq model with dual attention mechanisms to distinguish primary from secondary symptoms and map herb-symptom interactions using clinical data. Herb-Know [6] utilizes herb descriptions to model associations with symptoms and evaluates whether herb effects align with symptom descriptions, and other related work like TCMBERT [7], RoKEPG [8] also have shown their effectiveness on TCM prescription prediction.
@@ II-C Large Language Models in TCM
LLMs have advanced significantly in NLP tasks, with models like ChatGPT and GPT-4 [16] drawing attention, though OpenAI has not disclosed architectural details. Open-source LLMs, such as LLaMA [17] and ChatGLM [18], have spurred research interest, though they often struggle in specialized fields like medicine or finance. Supervised fine-tuning has become the standard to enhance these models with domain-specific knowledge. In TCM, tailored models have emerged. Bentsao [19] fine-tunes LLaMA using structured and unstructured knowledge from CMeKG [20]. Huatuo [21] incorporates ChatGPT data with Reinforcement Learning from Artificial Intelligence Feedback (RLAIF) and Proximal Policy Optimization (PPO). Zhongjing [22] applies continual pretraining with supervised fine-tuning on a Chinese medical dialogue dataset, integrating Reinforcement Learning from Human Feedback (RLHF). ShenNong [23] combines LoRA-based fine-tuning on instructional data and TCM knowledge, supported by over 110,000 instructions.
@@ II-D Parameter Efficient Fine-Tuning
Parameter Efficient Fine-Tuning (PEFT) utilizes a small amount of parameters to fine-tune a large language model effectively. Assuming there is a pretrained model $f_{\theta}(y|x)$, PEFT seeks to adjust a limited number of parameters, $\Delta\theta$, such that $|\theta|\gg|\Delta\theta|$. In contrast to conventional fine-tuning, which updates all parameters, denoted as $|\theta|=|\Delta\theta|$, and requires significant computational resources, PEFT selectively updates a limited number of learnable parameters to achieve results comparable to complete fine-tuning. In this work, we employ LoRA [12] to efficiently fine-tune LLMs on the target dataset.
@@ III-A1 Data collection
We collect outpatient medical record data from practical patient records generated by specialists in TCM hospital 111by Dr. Runshun Zhang in Guang’anmen Hospital, China over a span from 2008 to 2022. The prescriptions specifically focus on digestion system disorders.
@@ III-B TCM-FTP
Given a prescription dataset $\mathcal{P}_{train}$ with $\{s,\{h_{i},w_{i}\}^{i\in[k]}\}$, our primary goal is to use a language model to minimize the negative logarithm of the estimation of conditional probability:
$\min\limits_{\theta}\mathop{\mathbb{E}}\limits_{s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}}\sum\limits_{i}^{k}-\log P((h_{i},w_{i})|h_{<i},w_{<i};s;\theta),$ (1)
@@
We concatenate the herbs and dosages into one string to use the auto-regressive mechanism in LLMs and get the goal:
$\min\limits_{\theta}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}\\ y=\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]})\end{subarray}}\sum\limits_{j=1}^{|y|}-\log P(y_{j}|\{y_{<j}\};s;\theta),$ (2)
where $\{y_{<j}\}$ denotes the set of the tokens from index $1$ to index $j-1$ in $y$, and $|y|$ represents the number of tokens in $y$. $\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]})$ is the concatenated string.
@@ III-B1 Low-rank Adaptation
We employ the LoRA [12] to reduce computational resources. Towards $f_{\theta}$, we encounter two primary types of parametric functions within our model framework: linear and embedding functions, defined as:
${\mathcal{F}_{\text{linear}}}(x)={W_{l}}\cdot x,{\mathcal{F}_{\text{emb}}}(x)=\text{EMB}\left(x;{W_{e}}\right),$ (3)
where EMB$(x;\cdot)$ represents the embedding operator that selects the $x$-th column from the specified matrix, and $W_{l}$ and $W_{e}$ denote the parameters for the linear and embedding functions, respectively. For fine-tuning, we introduce updates via low-rank matrices $A$ and $B$ to modify the original functions:
$\begin{split}{\mathcal{F}_{\text{linear}}}{(x)^{\prime}}&={\mathcal{F}_{\text{linear}}}(x)+{A_{l}}\cdot{B_{l}}\cdot x,\\ {\mathcal{F}_{\text{emb}}}{(x)^{\prime}}&={\mathcal{F}_{\text{emb}}}(x)+{A_{e}}\cdot\text{EMB}\left(x;{B_{e}}\right),\end{split}$ (4)
where $A_{l}$, $B_{l}$, $A_{e}$, and $B_{e}$ are the learnable parameters. With the updates to the low-rank matrices represented by $\Delta\theta$, our goal is:
$\min\limits_{\Delta\theta}\mathop{\mathbb{E}}\limits_{\begin{subarray}{c}s,\{h_{i},w_{i}\}^{i\in[k]}\sim\mathcal{P}_{train}\\ y=\textsc{cat}(\{h_{i},w_{i}\}^{i\in[k]}\end{subarray}}\sum\limits_{j=1}^{|y|}-\log P(y_{j}|\{y_{<j}\};s;\theta+\Delta\theta).$ (5)
@@ III-C Baselines
To demonstrate the performance of our model, we compare it with several baselines, including TCM prescription prediction models (PTM [3], TCMPR [24], KDHR [25], and PresRecST [10]), and pre-trained language models (Mengzi (T5-base) [26], GPT-3.5 and GPT-4.0 [16]).